OpenStudy (anonymous):

choose what best completes the statement 2(5-3x)= 1/16 a.{8} b.{-3} c.{1/8} d.{3}

OpenStudy (anonymous):

it doesn't work?

OpenStudy (anonymous):

none, but the answer would be 477/32

OpenStudy (anonymous):

i mean 53/32

OpenStudy (anonymous):

hmmm thats weird

OpenStudy (anonymous):

it just doesn't work with a-d ^^ just solve the equation

OpenStudy (anonymous):

\[2^{(5-3x)}\] i think the first part may look like this instead

OpenStudy (anonymous):

well can't check this one with mental math xD

OpenStudy (anonymous):

it's d don't think i needed 6 min just started :P

OpenStudy (anonymous):

if the problem is 2 raised to the power (5−3x) = 1/16, then taking log[base 2] of both sides of equation LHS and RHS, gives us log[base 2] of 2 raised to the power (5−3x) = log[base 2]1/16, then the LHS is 5-3x and the right RHS is log[base 2](1/16). So equation is now 5-3x = -4, then 5+4 = 3x, and 9/3 = x. Thus x = 3...

OpenStudy (anonymous):

thanks guys